### v1

@@ Abstract
keywords:
[…]
{highlights}
Extend Harrison’s center theory to a set of multivariate polynomials.
Discover the connection between simultaneous direct sum decompositions and centers.
Provide an algorithm for simultaneously decomposing multivariate polynomials.
1 Introduction
1.1 Problem statement
Throughout, $\mathbb{k}$ is a field and $\mathbb{k}[x_{1},x_{2},\dots,x_{n}]$ is the polynomial ring in $n$ variables. Let $f_{i}(x)=f_{i}(x_{1},x_{2},\dots,x_{n})\in\mathbb{k}[x_{1},x_{2},\dots,x_{n}]$ for $1\leq i\leq m$. We are concerned about the problem of whether there exists a change of variables $x=Py$ such that
$f_{i}(x)=f_{i}(Py)=g_{i1}(y_{1},\ldots,y_{a_{1}})+g_{i2}(y_{a_{1}+1},\ldots,y_{a_{2}})+\cdots+g_{it}(y_{a_{t-1}+1},\ldots,y_{n})$ (1.1)
@@ 1.3 Methods and results
Let $f_{i}(x_{1},x_{2},\dots,x_{n})\in\mathbb{k}[x_{1},x_{2},\dots,x_{n}]$ for $1\leq i\leq m$ be as above. The center of this set of polynomials is defined as
@@ 1.4 Organization of the paper
2 The center of a set of multivariate polynomials
2.1 Some examples
@@
$Z(f_{1},f_{2})=\left\{\left(\begin{array}[]{cc}a&\frac{1}{2}(b-a)\\ \frac{3}{2}(b-a)&b\end{array}\right)\Big{|}a,b\in\mathbb{k}\right\}.$
@@ 2.2 Jordan algebras
$X\odot Y=Y\odot X,\quad(X^{2}\odot Y)\odot X=X^{2}\odot(Y\odot X).$ (2.1)
@@
In a unital Jordan subalgebra $J$ of $({\mathbb{k}^{n\times n}})^{+},$ if $e_{1},e_{2},\dots,e_{s}$ are a complete set of orthogonal idempotents, then we have
@@
2.3 Main properties of centers
@@
Let $Z(f_{1},f_{2},\ldots,f_{m})$ be the center of $f_{1},f_{2},\ldots,f_{m}\in\mathbb{k}[x_{1},x_{2},\dots,x_{n}]$.
@@
    If the coefficients of $f_{1},f_{2},\ldots,f_{m}$ are generic and at least one of the $f_{i}$ has degree not less than $3,$ then $(Z(f_{1},f_{2},\ldots,f_{m}),\odot)\cong\mathbb{k}$.
@@
for $1\leq i\leq m$. In other words, $(Z(f_{1},f_{2},\ldots,f_{m}),\odot)$ is closed under $\odot$ and hence a Jordan subalgebra of ${(\mathbb{k}^{n\times n})}^{+}$.
(2) Without loss of generality, we assume $f_{1}$ has at least two variables of degree $d\geq 3$. If $f_{1}$ has more than two variables or has degree $d>3$, then a generic $f_{1}=\sum_{l=0}^{d}f_{1}^{(l)}$ has nondegenerate degree $d$ homogeneous component $f_{1}^{(d)}$. By [10, Theorem 3.2], $Z(f_{1})\cong\mathbb{k}$ for a generic $f_{1}$ as it is clear by Definition 1.1 that $Z(f_{1})=\cap_{l=0}^{d}Z(f_{1}^{(l)}).$ If $f_{1}$ is a binary polynomial of degree $3,$ say
@@
where $A,B,C,D,E,F,G,H,I,K\in\mathbb{k},$ then the center $Z(f_{1})$ is the solution of the following linear equations
@@
So for generic $A,B,C,D,E,F,G,$ the previous linear equations have trivial solutions, namely, $Z(f_{1})$ consists of scalar matrices. That is, $Z(f_{1})\cong\mathbb{k}.$ It is obvious that $Z(f_{1},f_{2},\dots,f_{m})=\cap_{i=1}^{m}Z(f_{i})\cong\mathbb{k}$ for a general set of multivariate polynomials $f_{1},f_{2},\ldots,f_{m}$ with the prescribed conditions. ∎
@@
3 Simultaneous direct sum decompositions of polynomials
3.1 Center and simultaneous direct sum decompositions
@@
Suppose $f_{1},f_{2},\ldots,f_{m}\in\mathbb{k}[x_{1},x_{2},\dots,x_{n}].$ Then
@@
Conversely, suppose there exists a complete set of orthogonal idempotents $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ in $Z(f_{1},f_{2},\ldots,f_{m}).$ By Remark 2.4, we have $\epsilon_{i}^{2}=\epsilon_{i}$ for all $i$ and $\epsilon_{i}\epsilon_{j}=0$ for all $i\neq j$ as matrices. Since the $\epsilon_{i}$’s are diagonalizable matrices and they mutually commute, $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ are simultaneously diagoanalizable. Thus, there exists some invertible matrix $P\in\mathbb{k}^{n\times n}$ such that $P^{-1}\epsilon_{j}P=E_{j}$ as above and $\sum_{j=1}^{t}E_{j}=I_{n}$. Take the change of variables $x=Py$ and let $g_{i}$ and $G_{i}$ be as above. We now have $G_{i}E_{j}=(G_{i}E_{j})^{T}=E_{j}^{T}G_{i}^{T}=E_{j}G_{i}$ for $1\leq i\leq m$ and $1\leq j\leq t$. It follows that the $G_{i}$’s are quasi-diagonal and $g_{1},g_{2},\ldots,g_{m}$ can be written as sums of polynomials in disjoint sets of variables
@@
3.2 Algorithm for simultaneous direct sum decompositions
@@ 4 Some examples
$Z(f_{1},f_{2})=\left\{\left(\begin{array}[]{cc}a&\frac{1}{2}(b-a)\\ \frac{3}{2}(b-a)&b\end{array}\right)\Big{|}a,b\in\mathbb{k}\right\}$
@@
$Z(f_{1},f_{2})=\left\{\left(\begin{array}[]{cccc}a+b+c&a+2c&a&0\\ 0&b-c&c&0\\ 0&0&b&0\\ 0&0&0&b\end{array}\right)\Bigg{|}a,b,c\in\mathbb{k}\right\}.$
@@
$Z(f_{1})=\left\{\left(\begin{array}[]{ccc}a+3b&0&b\\ 0&a&0\\ 0&0&a\end{array}\right)\Bigg{|}a,b\in\mathbb{k}\right\}$
@@
$Z(f_{2})=\left\{\left(\begin{array}[]{ccc}a-\frac{1}{2}b&b&0\\ 0&a&0\\ 0&c&a\end{array}\right)\Bigg{|}a,b,c\in\mathbb{k}\right\}$
@@
$Z(f_{3})=\left\{\left(\begin{array}[]{ccc}a&0&0\\ e&d&b\\ \frac{3}{2}e&c&d-\frac{2}{3}c+\frac{3}{2}b\end{array}\right)\Bigg{|}a,b,c,d,e\in\mathbb{k}\right\}$
@@
$Z(f_{1},f_{2},f_{3})=\left\{\left(\begin{array}[]{ccc}a&0&0\\ 0&a&0\\ 0&0&a\end{array}\right)\Bigg{|}a\in\mathbb{k}\right\},$

### v2

@@ Abstract.
Key words and phrases:
multivariate polynomial, direct sum decomposition
2020 Mathematics Subject Classification:
15A69, 13P05
Supported by Key Program of Natural Science Foundation of Fujian Province (Grant no. 2024J02018) and National Natural Science Foundation of China (Grant no. 12371037).
1. Introduction
1.1. Problem statement
Throughout, $\mathbbm{k}$ is a field and $\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ is the polynomial ring in $n$ variables. Let $f_{i}(x)=f_{i}(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ for $1\leq i\leq m$. We are concerned about the problem of whether there exists a change of variables $x=Py$ such that
(1.1) $f_{i}(x)=f_{i}(Py)=g_{i1}(y_{1},\ldots,y_{a_{1}})+g_{i2}(y_{a_{1}+1},\ldots,y_{a_{2}})+\cdots+g_{it}(y_{a_{t-1}+1},\ldots,y_{n})$
@@ 1.3. Methods and results
Let $f_{i}(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ for $1\leq i\leq m$ be as above. The center of this set of polynomials is defined as
@@ 1.4. Organization of the paper
2. The center of a set of multivariate polynomials
2.1. Some examples
@@
$Z(f_{1},f_{2})=\left\{\left(\begin{array}[]{cc}a&\frac{1}{2}(b-a)\\ \frac{3}{2}(b-a)&b\end{array}\right)\Big{|}a,b\in\mathbbm{k}\right\}.$
@@ 2.2. Jordan algebras
(2.1) $X\odot Y=Y\odot X,\quad(X^{2}\odot Y)\odot X=X^{2}\odot(Y\odot X).$
@@
In a unital Jordan subalgebra $J$ of $({\mathbbm{k}^{n\times n}})^{+},$ if $e_{1},e_{2},\dots,e_{s}$ are a complete set of orthogonal idempotents, then we have
@@
2.3. Main properties of centers
@@
Let $Z(f_{1},f_{2},\ldots,f_{m})$ be the center of $f_{1},f_{2},\ldots,f_{m}\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$.
@@
    If the coefficients of $f_{1},f_{2},\ldots,f_{m}$ are generic and at least one of the $f_{i}$ has degree not less than $3,$ then $(Z(f_{1},f_{2},\ldots,f_{m}),\odot)\cong\mathbbm{k}$.
@@
for $1\leq i\leq m$. In other words, $(Z(f_{1},f_{2},\ldots,f_{m}),\odot)$ is closed under $\odot$ and hence a Jordan subalgebra of ${(\mathbbm{k}^{n\times n})}^{+}$.
(2) Without loss of generality, we assume $f_{1}$ has at least two variables of degree $d\geq 3$. If $f_{1}$ has more than two variables or has degree $d>3$, then a generic $f_{1}=\sum_{l=0}^{d}f_{1}^{(l)}$ has nondegenerate degree $d$ homogeneous component $f_{1}^{(d)}$. By [10, Theorem 3.2], $Z(f_{1})\cong\mathbbm{k}$ for a generic $f_{1}$ as it is clear by Definition 1.1 that $Z(f_{1})=\cap_{l=0}^{d}Z(f_{1}^{(l)}).$ If $f_{1}$ is a binary polynomial of degree $3,$ say
@@
where $A,B,C,D,E,F,G,H,I,K\in\mathbbm{k},$ then the center $Z(f_{1})$ is the solution of the following linear equations
@@
So for generic $A,B,C,D,E,F,G,$ the previous linear equations have trivial solutions, namely, $Z(f_{1})$ consists of scalar matrices. That is, $Z(f_{1})\cong\mathbbm{k}.$ It is obvious that $Z(f_{1},f_{2},\dots,f_{m})=\cap_{i=1}^{m}Z(f_{i})\cong\mathbbm{k}$ for a general set of multivariate polynomials $f_{1},f_{2},\ldots,f_{m}$ with the prescribed conditions. ∎
@@
3. Simultaneous direct sum decompositions of polynomials
3.1. Center and simultaneous direct sum decompositions
@@
Suppose $f_{1},f_{2},\ldots,f_{m}\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}].$ Then
@@
Conversely, suppose there exists a complete set of orthogonal idempotents $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ in $Z(f_{1},f_{2},\ldots,f_{m}).$ By Remark 2.4, we have $\epsilon_{i}^{2}=\epsilon_{i}$ for all $i$ and $\epsilon_{i}\epsilon_{j}=0$ for all $i\neq j$ as matrices. Since the $\epsilon_{i}$’s are diagonalizable matrices and they mutually commute, $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ are simultaneously diagoanalizable. Thus, there exists some invertible matrix $P\in\mathbbm{k}^{n\times n}$ such that $P^{-1}\epsilon_{j}P=E_{j}$ as above and $\sum_{j=1}^{t}E_{j}=I_{n}$. Take the change of variables $x=Py$ and let $g_{i}$ and $G_{i}$ be as above. We now have $G_{i}E_{j}=(G_{i}E_{j})^{T}=E_{j}^{T}G_{i}^{T}=E_{j}G_{i}$ for $1\leq i\leq m$ and $1\leq j\leq t$. It follows that the $G_{i}$’s are quasi-diagonal and $g_{1},g_{2},\ldots,g_{m}$ can be written as sums of polynomials in disjoint sets of variables
@@
3.2. Algorithm for simultaneous direct sum decompositions
@@ 4. Some examples
$Z(f_{1},f_{2})=\left\{\left(\begin{array}[]{cc}a&\frac{1}{2}(b-a)\\ \frac{3}{2}(b-a)&b\end{array}\right)\Big{|}a,b\in\mathbbm{k}\right\}$
@@
$Z(f_{1},f_{2})=\left\{\left(\begin{array}[]{cccc}a+b+c&a+2c&a&0\\ 0&b-c&c&0\\ 0&0&b&0\\ 0&0&0&b\end{array}\right)\Bigg{|}a,b,c\in\mathbbm{k}\right\}.$
@@
$Z(f_{1})=\left\{\left(\begin{array}[]{ccc}a+3b&0&b\\ 0&a&0\\ 0&0&a\end{array}\right)\Bigg{|}a,b\in\mathbbm{k}\right\}$
@@
$Z(f_{2})=\left\{\left(\begin{array}[]{ccc}a-\frac{1}{2}b&b&0\\ 0&a&0\\ 0&c&a\end{array}\right)\Bigg{|}a,b,c\in\mathbbm{k}\right\}$
@@
$Z(f_{3})=\left\{\left(\begin{array}[]{ccc}a&0&0\\ e&d&b\\ \frac{3}{2}e&c&d-\frac{2}{3}c+\frac{3}{2}b\end{array}\right)\Bigg{|}a,b,c,d,e\in\mathbbm{k}\right\}$
@@
$Z(f_{1},f_{2},f_{3})=\left\{\left(\begin{array}[]{ccc}a&0&0\\ 0&a&0\\ 0&0&a\end{array}\right)\Bigg{|}a\in\mathbbm{k}\right\},$